Q6) The wage rate of labor is Rs. 6 and price of capital is Rs. 2. The marginal product of labor is 16 while marginal product of capital is 4. Can a firm be operating at equilibrium?


1
Expert's answer
2021-10-31T18:30:40-0400

The equilibrium condition for the firm having optimal input combinations is that the marginal products of each input are proportional to the price of the inputs. If there are two inputs, labor (L) and capital (K) having marginal productivities as MPL and MPK respectively and the input prices of each input is w (wages) and r (the price of capital), then the formula for firm’s equilibrium is-

"\frac{MP_L}{w}=\frac{MP_K}{r}"


By substituting these values in the equilibrium condition, we get-

"\frac{16}{6}=\frac{4}{2}"

This shows that the given input combination is inefficient as the firm is able to produce more output when an extra unit of money is spent on labor than on capital. Therefore, the firm is not operating at equilibrium, and to work efficiently at equilibrium, the firm should substitute labor for capital.